OpenStudy (anonymous):

Solve the system of equations and enter the solution as an ordered pair. 4y = 3x + 5 6y = 3x + 21

OpenStudy (anonymous):

subtract 1st eqn from second one 6y = 3x + 21 _ 4y =3x + 5 ______________ 2y = 0 + 16 ______________ 2y=16 y=8 substitute y=8 to get x 4y = 3x + 5 3x+5=32 3x=27 x=9 answer is (9,8)